OpenStudy (anonymous):

Circle P has a radius of 8 units with center P at (2, -1).Which equation defines circle P? a. (x + 2)2 + (y – 1)2 = 64 b.(x – 2)2 + (y + 1)2 = 64 c.(x – 2)2 + (y + 1)2 = 8 d. (x + 2)2 + (y – 1)2 = 8

OpenStudy (anonymous):

The trick here is that if the center is at x = 2, the equation will have (x-2)^2

OpenStudy (anonymous):

\[(x - h)^2 + (y - k)^2 = r^2\]where r=radius h=x k=y

OpenStudy (anonymous):

And of course the other thing is that as Romeo says, you need r^2 on the rhs.

OpenStudy (anonymous):

So it's b or c.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no you can only have one answer.

OpenStudy (anonymous):

you square the radius and you get 64

OpenStudy (anonymous):

so it's c?

OpenStudy (anonymous):

what does c have in the end?

OpenStudy (anonymous):

8^2 is equal to 64.... but it's missing the ^2.

OpenStudy (anonymous):

If the radius is 8 then you will have 64 on the rhs.

OpenStudy (anonymous):

so we need a 64 in the end.

OpenStudy (anonymous):

we can say that c and d are not the answer because of this

OpenStudy (anonymous):

ok. thank you so much! Sometimes, it's hard to get some concepts trough my head. So, it's b

OpenStudy (anonymous):

Thanks!